OpenStudy (anonymous):

In a survey, three out of seven people named blue as their favorite color. One out of six named red. If 1,092 people were included in the survey, how many named neither blue nor red as their favorite color

OpenStudy (campbell_st):

the easy way to do this if find the blue people... 3/7 * 1092 = answer 1 find the red people 1/6 * 1092 = answer 2 Neither red nor blue = 1092 -(answer 1 + answer 2)

OpenStudy (anonymous):

ok thanks! :)